OpenStudy (anonymous):

Find y': y= (x + 1/x)(x - 1/x)

OpenStudy (anonymous):

Distribute the (x - 1/x) into the (x + 1/x) like before to get: (x + 1/x)x + (x + 1/x)(-1/x) = x^2 + 1 - 1 - 1/x^2 = x^2 - 1/x^2 Also equal to x^2 - x^-2. Can you find the derivative from here?

OpenStudy (anonymous):

what does the " ' " mean? (y')

OpenStudy (anonymous):

y' is read as "y prime" and is equal to the derivative of y.

OpenStudy (anonymous):

ah thanks again!